OpenStudy (anonymous):

use the four step process fo find f1(x) and then find f1(1), f1(2) and f1(3)....... f(x)= 8x/4+x

OpenStudy (harsimran_hs4):

what does f1(x) mean? do you mean f(1)?

OpenStudy (anonymous):

no it sort of looks like f ' (x), f ' (1) and so on

OpenStudy (harsimran_hs4):

ok then that means derivative f(x) = 8x/4 + x = 2x + x =3x have i written f(x) right or it is some thing else?

OpenStudy (anonymous):

Yes but this is telling me something about using this equation to find f '(x) = f(x+h)-f(x)/h

OpenStudy (harsimran_hs4):

if f(x) is 3x then then the derivative i.e f '(x) is 3 else you can use the formula which is that derivative (f '(x) ) is given by \[f \prime(x) = \lim_{h \rightarrow 0}\frac{ f(x+h) - f(x) }{ h } \] now can you find f '(x) ?

OpenStudy (anonymous):

f(x)=\[\frac{ 8x}{4+x}\]

OpenStudy (anonymous):

I have to find f ' (x) from that equation you gave me

OpenStudy (harsimran_hs4):

ok so so far we have been considering f(x) wrong..... let`s start again then..... can you tell me what is f(x+h)?

OpenStudy (anonymous):

I don't know...okay ive confused you! My problem is : use the four step process to find f ' (x) and then find f ' (1), f ' (2), f ' (3). f(x)=\[\frac{ 8x}{4+x}\]

OpenStudy (anonymous):

I was just assuming that in order to find f ' (x) we had to use the equation that you gave with the f(x+h)-f(x)/h

OpenStudy (harsimran_hs4):

in that process your 1st step still is f(x+h) do you know how to calculate it or not?

OpenStudy (harsimran_hs4):

www.math.wvu.edu/~hjlai/Teaching/Tip-Pdf/Tip1-5.pdf check this out .....it will help you out a bit in getting hold of method

OpenStudy (anonymous):

I dont know how to calculate that and I cant seem to find any examples to show me the way

OpenStudy (harsimran_hs4):

ok f(x) = 8x/(4+x) now to calculate f(x+h) all you nees is put x + h instead of x so f(x+h) = 8(x+h)/(4+(x+h ))

OpenStudy (harsimran_hs4):

2nd step : tell what is f(x+h) - f(x) i.e f(x+h) - f(x) = ?

OpenStudy (anonymous):

8x+8h/4x+4h=2x+2h ?

OpenStudy (harsimran_hs4):

no what i`ll do is now give you step by step solution and you try to understand each step (why this step?) ok

OpenStudy (harsimran_hs4):

f(x) = 8x/(4+x) f(x+h) = 8(x+h)/ (4 + (x+h)) clear till now?

OpenStudy (anonymous):

okay, thank you

OpenStudy (anonymous):

I understand the equation just not how to solve it, yes

OpenStudy (harsimran_hs4):

ok cool now f(x+h) - f(x) = 8(x+h)/ (4 + (x+h)) - 8x/(4+x) taking common denominator and simplifying we get f(x+h) - f(x) = 32h/(4+x)(4+x+h) clear till now?

OpenStudy (anonymous):

yes

OpenStudy (harsimran_hs4):

cool 3rd step was to take out common factor h from this expression so can you see that we have already obtained that in numerator 32h = 32(h) ? clear till now?

OpenStudy (anonymous):

yes, got it

OpenStudy (harsimran_hs4):

ok now putting the things in the above mentioned formula f '(x) = [ 32h/(4+x)(4+x+h) ] / h cancelling common factors 32/(4+x)(4+x+h) now taking the limit h tending to zero for the above expression we get f '(x) = 32/(4+x)(4+x) = 32/(4+x)^2

OpenStudy (harsimran_hs4):

final answer \[f \prime (x) = \frac{ 32 }{ (4+x)^{2} }\]

OpenStudy (anonymous):

alright, and that is the formula I use to find the other 3 answers?

OpenStudy (harsimran_hs4):

yes this is the general formula for all x expect -4 beacause it`s not in domain of the function

OpenStudy (harsimran_hs4):

also i appologize for the late reply we had a unexpected network failure....

OpenStudy (anonymous):

so f ' (1) = 1.28?

OpenStudy (harsimran_hs4):

yes!!

OpenStudy (anonymous):

Thats okay I appreciate you helping me out!

OpenStudy (harsimran_hs4):

my pleasure so finally you trough with the question or still some confusion?